OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7).

OpenStudy (austinl):

Do you know the formula to find slope?

OpenStudy (anonymous):

-.- wow dragonflame

OpenStudy (anonymous):

M=y2-y1/x2-x1

OpenStudy (austinl):

Correct, to clean it up a bit. \(\Large{m=\frac{y_2-y_1}{x_2-x_1}}\) And then you have two points. \((5, 1)\rightarrow(x_1,y_1)\) \((-4,7)\rightarrow(x_2,y_2)\) can you find slope for me?

OpenStudy (anonymous):

y=6 x=-9

OpenStudy (austinl):

\(\large{m=-\frac{6}{9}}\) Okay so now we have that. The basic form of a line is slope intercept is this, \(y=mx+b\) You put in the slope that you calculated. \(\large{y=-\frac{6}{9}x+b}\) Now, you need to take one of the points from above, and put them in for x and y to solve for b. Do you think you could do that?

OpenStudy (anonymous):

Yea one sec lemme do that

OpenStudy (anonymous):

\[y=-6/9*5+1\]

OpenStudy (austinl):

\(\large{1=-\frac{6}{9}(5)+b}\) Solve for b.

OpenStudy (anonymous):

Ah my bad sorry

OpenStudy (austinl):

No worries, it takes some time to get used to it :P

OpenStudy (anonymous):

30/9+1=b

OpenStudy (austinl):

Can you simplify that? 30/9+1?

OpenStudy (anonymous):

13/3

OpenStudy (austinl):

Correct, not plug that in for be in the basic formula with the m in as well.

OpenStudy (austinl):

now*

OpenStudy (anonymous):

What was the basic formula again?

OpenStudy (austinl):

\(y=mx+b\)

OpenStudy (anonymous):

13/3=-2/3x+b

OpenStudy (austinl):

No, plug it in for B

OpenStudy (anonymous):

y=-2/3x+13/3

OpenStudy (anonymous):

Math way helps

OpenStudy (austinl):

Correct Dragon, good job!

OpenStudy (anonymous):

Sure, but austin is helping me learn how to do it for future reference

OpenStudy (anonymous):

Thx

OpenStudy (austinl):

No problem, you have a good day/night/evening/whatevertimeitiswhereveryoulive I am gonna go study my calc 3, take care :)

OpenStudy (anonymous):

I will have a good 5:09 PM You 2!